# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@ Appendix A Appendix
$d_{\infty}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1},j}-h_{\theta_{2},j}\right\|_{\infty}\leq C_{\operatorname{Lip},\infty}\cdot\left(1+\left\|\theta_{1}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\left\|\theta_{2}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Using that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$, the polynomial growth can be bounded due to the embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$. Thus, it suffices to find an covering for $\Theta_{l}$ with respect to the $C^{\eta_{2}}$-norm. Indeed, by an application of Lemma A.9, we have
$N(\Theta_{l},\left\|\cdot\right\|_{C^{\eta_{2}(\mathcal{O},W)}},\rho)\leq\exp\left(c(\alpha,\eta_{2},d)\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\right).$
Now, for $\rho\in\mathbb{R}_{>0}$ fixed, let $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ be a covering of $\Theta_{l}$ w.r.t. $\left\|\cdot\right\|_{C^{\eta_{2}}(\mathcal{O},W)}$ with $N$ bounded by the line above. Thus, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
$d_{\infty}(\theta,\theta^{(i_{\circ})})\leq c(C_{\operatorname{Lip},\infty},\gamma_{2})\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{2}}\right)\cdot\rho.$
Thus, there exists a constant $c_{4}=c_{4}(\alpha,\eta_{2},d,d_{W})\in\mathbb{R}_{>0}$, such that
$\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\leq c_{4}\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta}}\text{ with}m_{2}:=c(C_{\operatorname{Lip},\infty},\gamma_{2})\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{2}}\right).$
We are now able to upper bound the corresponding entropy integrals. Indeed, as $2(\alpha+\kappa)>d$, we have
$\mathcal{J}_{2}(\mathcal{H}_{l,j})=\int_{0}^{4v_{l}}\sqrt{\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)}\mathrm{d}\rho\leq\sqrt{c_{3}}\cdot\int_{0}^{4v_{l}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{2(\alpha+\kappa)}}\mathrm{d}\rho=\sqrt{c_{3}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}},$
@@
$\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}}$ $=c_{1}^{\tau_{1}}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)^{\tau_{1}}\cdot 2^{l\tau_{1}}\cdot R^{\tau_{1}}\cdot\mathfrak{r}^{-\tau_{1}}\cdot v_{l}^{-1-\tau_{1}}$
@@
Evidently, the upper bound of $\mathcal{J}_{2}(\mathcal{H}_{l,j})$ can be as small as desired by choosing $c_{1}$ accordingly large. Analogously, as $\alpha-\eta>d$, we have
$\mathcal{J}_{\infty}(\mathcal{H}_{l,j})=\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\mathrm{d}\rho\leq c_{4}\cdot\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\mathrm{d}\rho\leq c(\alpha,\eta_{2},d,d_{W},C_{\mathscr{G},\operatorname{B}})\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{\alpha-\eta_{2}}}.$
Using Eq. C5 with similar arguments as above, one can verify analogously verify
@@
with some $c_{6}=c_{6}(\alpha,\eta_{1},\eta_{2},\gamma_{2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},c_{1})\in\mathbb{R}_{>0}$, which can be also as small as desired by choosing $c_{1}$ sufficiently large. Now in the setting of Lemma A.7, for arbitrary $\bar{c}\in\mathbb{R}_{>0}$, such that $x:=\bar{c}N2^{2l}R^{2}\in\mathbb{R}_{\geq 1}$, we have for some constant $L\in\mathbb{R}_{>0}$
$L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)\leq L\left(c_{5}\cdot v_{l}^{2}\sqrt{N}+v_{l}\sqrt{\bar{c}N2^{2l}R^{2}}+N^{-\frac{1}{2}}(c_{6}\cdot 2^{2l}R^{2}N+8C_{\mathscr{G},\operatorname{B}}\bar{c}N2^{2l}R^{2})\right)$
$\leq\frac{\sqrt{N}v_{l}^{2}}{96\cdot d_{V}}\left(c_{5}\cdot c_{1}^{-(1-\gamma_{1}\tau_{1})}\cdot d_{V}\cdot 96\cdot L+L\cdot 96\cdot d_{V}\frac{\sqrt{\bar{c}}}{\sqrt{c_{1}}}+c_{6}\cdot 96\cdot d_{V}\cdot L+\frac{8C_{\operatorname{Lip},2}\bar{c}}{c_{1}^{2}}\cdot 96\cdot d_{V}\cdot L\right)\leq\sqrt{N}\frac{v_{l}^{2}\sigma}{96\cdot d_{V}},$
which is attainable for $c_{1}=c_{1}(\alpha,\eta_{2},\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},d_{V},\bar{c},\sigma)\in\mathbb{R}_{>0}$ large enough. Hence an applycation of Lemma A.7 yields
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq\sqrt{N}\frac{v_{l}^{2}\sigma}{96d_{V}}\right)$ $\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)\right)$
$\leq 2\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
@@
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{v_{l}^{2}}{96}\right)\leq 2d_{V}\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
@@
$T_{2}(\theta):=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\}=:\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}h_{\theta}(Z_{i}),$
we need to controll
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{\left|\frac{1}{\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}(h_{\theta}(Z_{i})-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[h_{\theta}(Z_{i})])\right|\right\}\geq\sqrt{N}\frac{2^{2l}}{48}R^{2}c_{1}^{2}\sigma^{2}\right\}\right).$
@@
$\mathcal{H}_{l}:=\left\{h_{\theta}:=\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\tilde{\theta})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta)(\cdot)\right|_{V}^{2}:\mathcal{Z}\to\mathbb{R}\,:\,\theta\in\Theta_{l}\right\},$
and verify the assumptions of Lemma A.7. Indeed, note that for each $\theta\in\Theta_{l}$, we have
$\sup_{z\in\mathcal{Z}}\left\{\left|h_{\theta}(z)\right|\right\}$ $=\sup_{z\in\mathcal{Z}}\left\{\left|\left|\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\tilde{\theta})(z)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\theta)(z)\right|_{V}^{2}\right|\right\}\leq 8C_{\mathscr{G},\operatorname{B}}$
@@
$\mathbb{E}^{Z}[h_{\theta}^{2}(Z)]$ $=\mathbb{E}^{Z}\left[\left|\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\tilde{\theta})(Z)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\theta)(Z)\right|_{V}^{2}\right|^{2}\right]\leq 16C_{\mathscr{G},\operatorname{B}}^{2}\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta)\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq 16C_{\mathscr{G},\operatorname{B}}^{2}v_{l}^{2},$
where we have applied C2) from 2.3 and used the notation as before. In order to bound the entropy integrals, observe that for any $\theta_{1},\theta_{2}\in\Theta_{l}$
$d_{2}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1}}-h_{\theta_{2}}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}=\left\|\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)\right|_{V}^{2}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}\leq 4C_{\mathscr{G},\operatorname{B}}\cdot\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)},$
@@
$d_{\infty}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1}}-h_{\theta_{2}}\right\|_{\infty}=\left\|\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)\right|_{V}^{2}\right\|_{\infty}$ $\leq 4C_{\mathscr{G},\operatorname{B}}\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\infty}.$
@@
$\log N\left(\mathcal{H}_{l},d_{2},\rho\right)\text{ and }\log N\left(\mathcal{H}_{l},d_{\infty},\rho\right),$
as already seen for the process $T_{1}$, by a change of constant. Thus, following the argumentation as before, we can choose $c_{1}:=(\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},d_{V},\bar{c},\sigma)\in\mathbb{R}_{>0}$ sufficiently large to obtain
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right)\leq 2\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
@@
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{7}\cdot\sum_{l\in\mathbb{N}}\exp(-\bar{c}N2^{2l}R^{2})\leq c_{7}\exp(-\bar{c}NR^{2}),$
@@
Let $\mathfrak{r}\in\mathbb{R}_{>0}$ satisfy Eq. C6 and fix some $\bar{c}\in\mathbb{R}_{\geq 1}$. We define the map $T:(\mathfrak{r}^{2},\infty)\to\mathbb{R}_{>0},\,t\mapsto c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+t)$ and set $\mathfrak{m}:=c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+\mathfrak{r}^{2})$, with $c_{1}$ from the first part of the theorem. Applying the transformation theorem, we obtain
@@
Now observe that $t\in(\mathfrak{r}^{2},\infty)$ satisfies $\sqrt{t}\geq\mathfrak{r}$ and hence also Eq. C4 and Eq. C5 by an application of Eq. C6. Thus, by an application of the first part of the theorem
$\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})]$ $\leq\mathfrak{m}+c_{1}\cdot c(d_{V})\cdot\int_{(\mathfrak{r}^{2},\infty)}\exp(-\bar{c}Nt)\mathrm{d}\mathcal{L}(t)\leq\mathfrak{m}+\frac{c_{1}\cdot c(d_{V})}{\bar{c}N},$
where we have chosen $c_{1}=c_{1}(\bar{c},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},\sigma)$ sufficiently large. This shows the claim.
Proof of 7 (Corollary 2.7).

### v3

@@ 4.1 Proofs of section 2
$d_{\infty}(\theta_{1},\theta_{2})=\|h_{\theta_{1},j}-h_{\theta_{2},j}\|_{\infty}\leq C_{\operatorname{Lip},\infty}\left(1+\|\theta_{1}\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\|\theta_{2}\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\|\theta_{1}-\theta_{2}\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Using that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$, the polynomial growth can be bounded due to the continuous embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$. Hence, it suffices to find a covering for $\Theta_{l}$ with respect to the $C^{\eta_{2}}$-norm. By an application of lemma 4.9 ii), we have
$N(\Theta_{l},\|\cdot\|_{C^{\eta_{2}(\mathcal{O},W)}},\rho)\leq\exp\left(c(\alpha,\eta_{2},d)\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\right).$
As a conseqence, for every $\rho\in\mathbb{R}_{>0}$, there exists a $\rho$-covering $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ of $\Theta_{l}$ with respect to $\|\cdot\|_{C^{\eta_{2}(\mathcal{O},W)}}$ with $N=N(\Theta_{l},\|\cdot\|_{C^{\eta_{2}(\mathcal{O},W)}},\rho)$. Thus, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
$d_{\infty}(\theta,\theta^{(i_{\circ})})\leq{\color[rgb]{0,0,0}\|h_{\theta,j}-h_{\theta^{(i_{\circ})},j}\|_{\infty}}\leq c(C_{\operatorname{Lip},\infty},\gamma_{2})\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{2}}\right)\cdot\rho{\color[rgb]{0,0,0}=:m_{2}\cdot\rho}$
and we deduce
$\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\leq c_{4}\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta}}$
for some constant $c_{4}=c_{4}(\alpha,\eta_{2},d,d_{W})\in\mathbb{R}_{>0}$. We are now able to upper bound the corresponding entropy integrals. Indeed, as $2(\alpha+\kappa)>d$, we have
$\mathcal{J}_{2}(\mathcal{H}_{l,j})$ $=\int_{0}^{4v_{l}}\sqrt{\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)}\mathrm{d}\rho$
$\leq\sqrt{c_{3}}\cdot\int_{0}^{4v_{l}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{2(\alpha+\kappa)}}\mathrm{d}\rho=\sqrt{c_{3}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}},$
@@
$\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}}\hskip-4.30554pt$ $=c_{1}^{\tau_{1}}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)^{\tau_{1}}\cdot 2^{l\tau_{1}}\cdot R^{\tau_{1}}\cdot\mathfrak{r}^{-\tau_{1}}\cdot v_{l}^{-1-\tau_{1}}$
@@
Analogously, as $\alpha-\eta_{2}>d$, we have
$\mathcal{J}_{\infty}(\mathcal{H}_{l,j})$ $=\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\mathrm{d}\rho$
$\leq c_{4}\cdot\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\mathrm{d}\rho\leq c(\alpha,\eta_{2},d,d_{W},C_{\mathscr{G},\operatorname{B}})\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{\alpha-\eta_{2}}}.$
Using eq. C5 with similar arguments as above, one can verify analogously
@@
with some $c_{6}=c_{6}(\alpha,\eta_{1},\eta_{2},\gamma_{2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},c_{1})\in\mathbb{R}_{>0}$, which can be also as small as desired by choosing $c_{1}$ sufficiently large. Now in the setting of lemma 4.7, for arbitrary $\bar{c}\in\mathbb{R}_{\geq 1}$, such that $x:=\bar{c}2^{2l}NR^{2}\in\mathbb{R}_{\geq 1}$, we have for some universal constant $L\in\mathbb{R}_{>0}$
$L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)$
$\hskip 28.45274pt\leq L\left(c_{5}c_{1}^{-(1-\gamma_{1}\tau_{1})}v_{l}^{2}\sqrt{N}+v_{l}\sqrt{\bar{c}}\sqrt{N}2^{l}R+\frac{1}{\sqrt{N}}(c_{6}v_{l}^{2}N+8\bar{c}N2^{2l}R^{2}C_{\mathscr{G},\operatorname{B}})\right)$
$\hskip 28.45274pt\leq\sqrt{N}\frac{v_{l}^{2}\sigma}{96\cdot d_{V}}\left(96Ld_{V}\sigma^{-1}c_{5}c_{1}^{-(1-\gamma_{1}\tau_{1})}+96Ld_{V}\frac{\sqrt{\bar{c}}}{\sigma c_{1}}+96Ld_{V}c_{6}+768Ld_{V}C_{\mathscr{G},\operatorname{B}}\frac{\bar{c}}{c_{1}^{2}}\right)$
$\hskip 28.45274pt\leq\sqrt{N}\frac{v_{l}^{2}\sigma}{96\cdot d_{V}},$
where in the last step follows after choosing $c_{1}=c_{1}(\texttt{P})\in\mathbb{R}_{>0}$ sufficiently large. Hence an application of lemma 4.7 yields
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq\sqrt{N}\frac{v_{l}^{2}\sigma}{96d_{V}}\right)$
$\hskip 42.67912pt\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)\right)$
$\hskip 42.67912pt\leq 2\exp\left(-\bar{c}2^{2l}NR^{2}\right).$
@@
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{v_{l}^{2}}{96}\right)\leq 2d_{V}\exp\left(-\bar{c}2^{2l}NR^{2}\right).$
@@
$T_{2}(\theta)=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}h_{\theta}(Z_{i})$
with $h_{\theta}(\cdot)=|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\tilde{\theta})(\cdot)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta)(\cdot)|_{V}^{2}$, we need to control
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{\left|\frac{1}{\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}(h_{\theta}(Z_{i})-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[h_{\theta}(Z_{i})])\right|\right\}\geq\sqrt{N}\frac{2^{2l}}{48}R^{2}c_{1}^{2}\sigma^{2}\right).$
@@
$\mathcal{H}_{l}:=\left\{h_{\theta}:=|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\tilde{\theta})(\cdot)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta)(\cdot)|_{V}^{2}:\mathcal{Z}\to\mathbb{R}\,:\,\theta\in\Theta_{l}\right\},$
and verify the assumptions of lemma 4.7. Indeed, note that for each $\theta\in\Theta_{l}$, we have
$\sup_{z\in\mathcal{Z}}\left\{\left|h_{\theta}(z)\right|\right\}$ $=\sup_{z\in\mathcal{Z}}\left\{\left||\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\tilde{\theta})(z)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\theta)(z)|_{V}^{2}\right|\right\}\leq 8C_{\mathscr{G},\operatorname{B}}$
@@
$\mathbb{E}^{Z}[h_{\theta}^{2}(Z)]$ $=\mathbb{E}^{Z}\left[\left||\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\tilde{\theta})(Z)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\theta)(Z)|_{V}^{2}\right|^{2}\right]$
$\leq 16C_{\mathscr{G},\operatorname{B}}^{2}\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta)\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq 16C_{\mathscr{G},\operatorname{B}}^{2}v_{l}^{2},$
with $v_{l}$ defined in eq. 24, where we have applied C2) from 2.3. In order to bound the entropy integrals, observe that for any $\theta_{1},\theta_{2}\in\Theta_{l}$
$d_{2}(\theta_{1},\theta_{2})$ $=\|h_{\theta_{1}}-h_{\theta_{2}}\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}=\||\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)|_{V}^{2}\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$
$\leq 4C_{\mathscr{G},\operatorname{B}}\cdot\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)},$
@@
$d_{\infty}(\theta_{1},\theta_{2})$ $=\|h_{\theta_{1}}-h_{\theta_{2}}\|_{\infty}=\||\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)|_{V}^{2}\|_{\infty}$
$\leq 4C_{\mathscr{G},\operatorname{B}}\cdot\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\infty}.$
@@
$\log N\left(\mathcal{H}_{l},d_{2},\rho\right)\text{ and }\log N\left(\mathcal{H}_{l},d_{\infty},\rho\right)$
as already derived before (up to a change of constants). Thus, following the same argumentation as above, we can choose $c_{1}=c_{1}(\texttt{P},d_{V})\in\mathbb{R}_{>0}$ sufficiently large to obtain
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right)\leq 2\exp\left(-\bar{c}2^{2l}NR^{2}\right).$
@@
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{7}\cdot\sum_{l\in\mathbb{N}}\exp(-\bar{c}2^{2l}NR^{2})\leq c_{7}\exp(-\bar{c}NR^{2}),$
@@
Let $\mathfrak{r}\in\mathbb{R}_{>0}$ satisfy eq. C6 and fix some $\bar{c}\in\mathbb{R}_{\geq 1}$. We define the map $T:(\mathfrak{r}^{2},\infty)\to\mathbb{R}_{>0},\,t\mapsto c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+t)$ and set $\mathfrak{m}:=c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+\mathfrak{r}^{2})$, with $c_{1}\in\mathbb{R}_{>0}$ choosen sufficiently large according to i). Applying the transformation theorem, we obtain
@@
Now observe that $t\in(\mathfrak{r}^{2},\infty)$ satisfies $\sqrt{t}\geq\mathfrak{r}$ and hence also eq. C4 and eq. C5 by an application of eq. C6. Thus, utilizing the concentration inequality yields
$\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})]$ $\leq\mathfrak{m}+c_{1}\cdot c(d_{V})\cdot\int_{(\mathfrak{r}^{2},\infty)}\exp(-\bar{c}Nt)\mathrm{d}\mathcal{L}(t)\leq\mathfrak{m}+\frac{c_{1}\cdot c(d_{V})}{\bar{c}N}.$
The claim then follows by taking the infimum over all admissible $\tilde{\theta}$. ∎
Proof of corollary 2.7.